OpenStudy (bohotness):

The width of a rectangle is 21 centimeters. If the perimeter of the rectangle is 102 centimeters, what is the length? 30 cm 81 cm 5 cm 60 cm

OpenStudy (bohotness):

a//?

jagr2713 (jagr2713):

l=P/2﹣w=102/2﹣21= what do you get

OpenStudy (bohotness):

30?

jagr2713 (jagr2713):

Yes great job :D